OpenStudy (anonymous):

NEED HELP! Find an equation in slope-intercept form of the line that has slope 5 and passes through point A(4,-5).

OpenStudy (anonymous):

A. y=5x+25 B. y=-5x-25 C. y=25x-5 D. y=5x-25

Nnesha (nnesha):

y=mx+b is a slope intercept form where m is slope and b is y-intercept so substitute m for 5 and then replace x and y with order pair (4,-5) solve for b

Nnesha (nnesha):

any question ?